Tohban Report 2013-08-07

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 31 July 2013
End Date: 07 August 2013
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: Mitsuo Oka
List all reports


Solar Activity

Currently there are 5 active regions on the disk, none of which are producing much in the way of flares. This extremely low level of activity is expected to continue.

How many GOES flares occurred?

 Flares above B, C, M, X class were     11     1     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      9     1     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      4     0     0     0

There were RHESSI flares/GOES flares 89 / 12 over the time range 31-Jul-13 07-Aug-13

Memory Management

No issues.

Spacecraft Status

The cold plate temperatures are dropping a little, as expected due to the increased day length.

Data Gaps

The large gaps from the previous week (July 29th/30th) due to the outage at Wallops appear to have been filled in (data replayed).

There is currently a gap from yesterday (August 6th) due to a temporary mishap at Wallops but we'll fill in that data soon.

Detector issues

The somewhat high slow rates in D7 continued this week. No action was taken. David reports that this "noise" is not affecting D7's spectroscopy, although the calibration has changed slightly.

There were isolated arcs in D2 on August 4th at 10:26 UT and 10:41 UT. (There may be hints in the other detectors, or it could be glitches.)

Other notes

Decimation Normal/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Enabled, during entire eclipse
Night time data (rears) Enabled, during entire eclipse
Require extra passes? No?
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? Detector 7 increased noise continued; Isolated small spikes in D2.
Personal tools